The molecular mass of penicillin G (an antibiotic) is 334. If you received a 1.9 ml injection of penicillin that contained 0.6 g of penicillin, what would be the concentration of the penicillin in the syringe?
1
Expert's answer
2015-09-01T03:05:09-0400
First determine the number of moles of penicillin G: n = 0.6g/334 = 0.0018 mol.
Then we can find the concentration: c = n/V = 0.0018mol/0.0019L = 0.95M.